# A Survey on Web Testing: On the Rise of AI and Applications in Industry

### v1

@@ Abstract
Web application testing is an essential practice to ensure the reliability, security, and performance of web systems in an increasingly digital world. This paper presents a systematic literature survey focusing on web testing methodologies, tools, and trends from 2014 to 2024. By analyzing $214214$ research papers, the survey identifies key trends, demographics, contributions, tools, challenges, and innovations in this domain. In addition, the survey analyzes the experimental setups adopted by the studies, including the number of participants involved and the outcomes of the experiments. Our results show that web testing research has been highly active, with ICST as the leading venue. Most studies focus on novel techniques, emphasizing automation in black-box testing. Selenium is the most widely used tool, while industrial adoption and human studies remain comparatively limited. The findings provide a detailed overview of trends, advancements, and challenges in web testing research, the evolution of automated testing methods, the role of artificial intelligence in test case generation, and gaps in current research. Special attention was given to the level of collaboration and engagement with the industry. A positive trend in using industrial systems is observed, though many tools lack open-source availability.
Keywords: web testing, GUI testing, front-end, survey, SBST, AI, fuzzing
@@ 1 Introduction
In this survey, we aim to identify key trends, advancements, and challenges of the web testing domain over the past decade. To achieve a comprehensive understanding of this field, we have analyzed $214214$ papers covering the period 2014-2024. Additionally, we have formulated 14 research questions, that attempt to explore and analyze the advancements, challenges, and trends in web testing since 2014.
The article is organized as follows. Section 2 introduces important background information needed to better understand the rest of this article. Section 3 discusses related work. Section 4 describes the paper selection process. In Section 5 we answer and discuss our research questions. Threats to validity are discussed in Section 6. Section 7 provides the conclusion of the article.
@@ 3 Related Work
There are three recent peer-reviewed surveys about web testing that are close to our work [94, 161, 19], analyzing from 26 to 98 (but this latter going back to 2008) articles, considering a similar time period to our 2014-2024 (see Table 1). By analyzing $214214$ articles, we can provide a more thorough view of the last decade in web testing research. Furthermore, with 14 research questions, we do analyze several important aspects of web testing that were not discussed before. For example, considering the current success stories of artificial intelligence techniques in many different aspects of science and engineering, we make sure to analyze how they have been used in web testing (as we will discuss in Section 5.13). Also, our survey focus only on testing. Other aspects like “automated repair” of code [146] and “fault localization” [171] (which were included for example in [19]) are not analyzed here in our survey, as not in scope.
A non-peer-reviewed survey [131], published on arXiv in 2024, covers similar topics compared to our survey. In [131] it is stated that 314 articles are reviewed, but only 199 are cited. So we cannot check if in our survey we missed any, or those other articles cover different aspects of web applications that we do not consider in our survey. Furthermore, with only six research questions [131], there are several important research questions that we study in our survey that are not covered in [131].
Compared to existing surveys in the literature, our survey provides the important contribution of a more in-depth analysis of the last 10 years of research in web frontend testing.
@@ 4.1 Research questions
  • RQ1: How many scientific articles were published each year between 2014-2024?
@@
  • RQ3: What kind of scientific contributions do these articles provide?
  • RQ4: What are the main scientific topics of focus in these articles?
@@
  • RQ6: Which tools are introduced or improved?
  • RQ7: How often tools are open-source?
  • RQ8: In algorithm experiments, how often are proposed tools compared with other tools?
  • RQ9: In human studies, what is the number of participants involved in the experiements?
  • RQ10: In case of algorithm experiment, how many faults are found?
  • RQ11: Which test generation technique is mostly used?
@@ 4.2 Databases and Search Queries
IEEE ( (”All Metadata”: ”web NEAR/2 test*”) OR (”All Metadata”: ”web NEAR/2 fuzz*”) OR (”All Metadata”: ”UI NEAR/2 test*”) OR (”All Metadata”: ”UI NEAR/2 fuzz*”) OR (”All Metadata”: ”GUI NEAR/2 test*”) OR (”All Metadata”: ”GUI NEAR/2 fuzz*”) OR ((”Document Title”: ”GUI” ) AND ((”All Metadata”: ”test” )OR (”All Metadata”:testing)) AND (”All Metadata”: ”web”)) OR ((”Document Title”: test* OR ”Document Title”: fuzz*) AND( (”Document Title”: ”web app*”))) ) AND NOT((”All Metadata”: fuzzy)) AND (NOT (”Document Title”:mobile OR ”Document Title”: android OR ”Document Title”: iOS OR ”Document Title”:desktop) ) 160
ACM Title: (test*) AND (NOT (Title:(mobile) OR Title: (android)) ) AND ((Abstract:(GUI) OR Abstract:(front-end) OR Abstract:(frontend) OR Abstract:(front end) ) AND (Abstract: (test*) OR Abstract:(fuzz*)) AND (Abstract: (Web)) ) 58
ScienceDirect Title, abstract, keywords: (”test” OR ”testing” OR ”fuzz” OR ”fuzzing”) AND (”user interface” OR ”UI” OR ”GUI”) AND (”web application” OR ”web”) 57
Wiley ”test* OR fuzz*” in Abstract and ”GUI OR frontend OR front-end OR front end OR user interface OR UI” in Abstract and ”test*” in Title 13
WebOfScience (((TI=(”test*” OR ”testing*” OR ”fuzz*” OR ”fuzzing*”)) AND AB=(”test*” OR ”testing*” OR ”fuzz*” OR ”fuzzing*”)) AND AB=(”web”)) AND AB=(”GUI” OR ”graphical user interface” OR ”UI” OR ”user interface” OR ”front end” OR ”front-end” OR ”frontend”) 61
SpringerLink (”test” OR ”testing” OR ”fuzz” OR ”fuzzing”) AND (”web application”) AND (”GUI” OR ”graphical user interface” OR ”UI” OR ”user interface” OR ”front end” OR ”front-end” OR ”frontend”) 236
MIT Title containing these search terms testing OR test OR fuzz OR fuzzing AND Abstract/description containing these search terms ”test” OR ”testing” OR ”fuzz” OR ”fuzzing” AND Abstract/description containing these search terms ”GUI” OR ”front end” OR ”front-end” OR ”front end” OR ”UI” or ”user interface” AND Abstract/description containing these search terms ”web application” 40
Total 625
Table  2 shows the queries we run in every database.
@@ 4.3 Paper Selection Criteria
Our initial search, conducted in 17-24/04/2023 resulted in 625 papers. This large number can be attributed to the fact that our main keywords are quite common and appear even in papers that might not be related to web testing. We decided to eliminate the papers that did not fit the criteria below.
@@
The first two authors went over all the papers independently, decided on the list of papers to be included, and then compared their respective lists. In case of any discrepency, the issue was solved by the fourth author. After checking all the papers retrieved initally from the databases, we ended up with $145145$ papers. These papers were reviewed again from the authors, and another $2121$ papers were excluded.
The second phase in paper collection involves snowballing. We followed the guidelines described in  [239]. We conducted forward and backward snowballing, and we ended up adding $9090$ new papers. Our final list amounts to $214214$ papers.
@@ 4.4 Data Extraction
The study comprises an extensive collection of papers. To manage and analyze the data, a structured Google spreadsheet was created. This document contains the list of selected papers alongside the extracted data, which is organized to address the research questions outlined in Section  4.1. Given the extensive collection of papers analyzed in this study, the dataset was divided between the first two authors for initial review and data extraction. Upon completion of the spreadsheet containing all extracted data, the results were collaboratively reviewed and thoroughly cross-verified by the remaining two authors to ensure accuracy and consistency.
@@ 5 Empirical Study
5.1 RQ1 - How many scientific articles were published each year between 2014-2024?
@@
Figure 1 illustrates the publication rate of papers on web testing from January 2014 to April 2024. The publication rate has remained relatively stable, fluctuating between 15 and 30 papers per year, with the peak reached in 2021. The lower number of publications in 2024 can be attributed to the fact that the paper selection process began in April 2024.
RQ1: The field of research on Web Testing has been significantly active in the last decade, with no signs of any slow down.
@@ 5.2 RQ2 - Which are the most common venues where these articles are published?
Table 3: RQ2 Results
Venue Nr.Papers Venue.Type
Other conferences 51 conference
Other journals 24 Journal
ICST 17 conference
FSE 11 conference
ICSE 10 conference
ICSTW 9 workshop
JSS 8 journal
STVR 7 journal
QUATIC 6 conference
SQJ 6 journal
TSE 6 journal
EMSE 5 journal
IST 5 journal
ASE 4 conference
AST 4 conference; workshop
ISSTA 4 conference
ICSME 3 conference
ICWE 3 conference
¡Programming¿ 2 conference
AUSE 2 journal
COMPSAC 2 conference
ESEM 2 conference
ICEBE 2 conference
IEEE Access 2 journal
IJSEKE 2 journal
ISMSIT 2 conference
ISSRE 2 conference
JSEP 2 journal
SoSyM 2 journal
SSBSE 2 conference
Table 3 shows all the venues in which the different surveyed articles have been published. These venues are sorted based on the number of published articles. For each venue, it is indicated if it is a journal, a conference or a workshop. Due to the large number of different venues (more than 100), we only report the acronym of the venues in which at lest two articles have been published. All others are grouped together into the “Other conferences” and “Other journals” categories.
@@
Regarding journals, the most commons are JSS (Journal of Systems and Software), STVR (Software Testing, Verification and Reliability), SQJ (Software Quality Journal), TSE (IEEE Transactions on Software Engineering), EMSE (Empirical Software Engineering) and IST (Information and Software Technology). With the notable exception of the missing TOSEM (ACM Transactions on Software Engineering and Methodology), these are arguably the most important journals focused on software engineering and software testing.
@@
RQ2: ICST is the most common venue where to find the latest advances in Web Testing research, followed by FSE and ICSE. The top journals in software engineering all share a similar amount of articles on Web Testing, with the notable exception of TOSEM.
5.3 RQ3 - What kind of scientific contributions do these articles provide?
New Tool or Technique (NToT):
Algorithm Experiment (AE):
Human Experiment (HE):
Position Paper (PP):
Experience Report (ER):
Analysis of Open-Source Projects (OSA):
Survey and Interviews (survey):
Table 4: List of papers for every contribution
Type Nr.Papers List.papers
NToT;AE 134 [160] [162] [137] [250] [46] [204] [248] [105] [192] [23] [96] [178] [67] [31] [187] [149] [243] [41] [138] [66] [251] [35] [158] [245] [50] [72] [36] [100] [144] [163] [235] [218] [111] [102] [234] [142] [113] [193] [134] [157] [240] [203] [247] [221] [185] [244] [151] [88] [155] [16] [202] [132] [86] [190] [198] [106] [173] [169] [103] [246] [191] [125] [61] [101] [56] [10] [179] [167] [172] [4] [210] [143] [60] [24] [26] [129] [90] [233] [98] [124] [207] [99] [227] [62] [219] [156] [183] [126] [216] [154] [215] [135] [97] [37] [205] [27] [229] [206] [174] [237] [92] [127] [123] [25] [213] [109] [13] [29] [189] [175] [42] [77] [141] [38] [107] [110] [7] [165] [220] [20] [47] [252] [166] [119] [57] [236] [39] [249] [14] [211] [65] [130] [64] [89]
$P$ 15 [214] [228] [73] [182] [231] [194] [199] [164] [196] [91] [212] [209] [69] [112] [201]
NToT 13 [32] [170] [44] [232] [114] [136] [49] [54] [79] [30] [85] [230] [43]
NToT;HE 11 [140] [68] [177] [74] [128] [95] [225] [184] [159] [150] [71]
HE 10 [195] [45] [17] [34] [28] [121] [115] [120] [117] [22]
ER 9 [33] [83] [52] [48] [15] [70] [53] [80] [176]
AE 8 [186] [18] [6] [108] [51] [208] [122] [147]
OSA 5 [93] [40] [145] [181] [200]
NToT;AE;HE 4 [63] [168] [139] [217]
survey 2 [118] [76]
AE;OSA 1 [87]
NToT;survey 1 [78]
NToT;AE;OSA 1 [242]
Table 5: Contributions
Contribution Nr.Papers
New Tool or Technique 164
Algorithm Experiment 148
Human Experiment 25
Position Paper 15
Experience Report 9
Analysis of Open-Source Projects 7
Survey and Interviews 3
RQ3: most of the research work on Web Testing still focus on designing novel techniques and tools, which are then evaluated “in the lab”. Their application in industry, and the study of their use among human subjects, have received significantly less attention.
5.4 RQ4 - What are the main scientific topics of focus in these articles?

### v2

@@ Abstract
Web application testing is an essential practice to ensure the reliability, security, and performance of web systems in an increasingly digital world. This paper presents a systematic literature survey focusing on web testing methodologies, tools, and trends from 2014 to 2025. By analyzing $259$ research papers, the survey identifies key trends, demographics, contributions, tools, challenges, and innovations in this domain. In addition, the survey analyzes the experimental setups adopted by the studies, including the number of participants involved and the outcomes of the experiments. Our results show that web testing research has been highly active, with ICST as the leading venue. Most studies focus on novel techniques, emphasizing automation in black-box testing. Selenium is the most widely used tool, while industrial adoption and human studies remain comparatively limited. The findings provide a detailed overview of trends, advancements, and challenges in web testing research, the evolution of automated testing methods, the role of artificial intelligence in test case generation, and gaps in current research. Special attention was given to the level of collaboration and engagement with the industry. A positive trend in using industrial systems is observed, though many tools lack open-source availability.
keywords:
Web Testing , GUI Testing , front-end , survey , SBST , AI , fuzzing
journal: Journal of Systems and Software
[…]
[label1]organization=Kristiania University of Applied Sciences, addressline=Kirkegata 24-26, city=Oslo, postcode=0153, country=Norway
[…]
[label2]organization=Glasgow Caledonian University, addressline=Cowcaddens Road, city=Glasgow, postcode=G4 0BA, state=Scotland, country=UK
{highlights}
Analyze Web Testing methodologies, tools and trends on the last decade (2014-2025)
Analyze the rate of collaboration with industry.
Analyze the role of AI in web testing.
@@ 1 Introduction
In this survey, we aim to identify key trends, advancements, and challenges of the web testing domain over the past decade. To achieve a comprehensive understanding of this field, we have analyzed $259$ papers covering the period 2014-2025. Additionally, we have formulated 14 research questions, that attempt to explore and analyze the advancements, challenges, and trends in web testing since 2014.
The article is organized as follows. Section 2 introduces important background information needed to better understand the rest of this article. Section 3 discusses related work. Section 4 describes the paper selection process. In Section 5 we answer and discuss our research questions. Our reflections and insight from answering these research questions are presented in Section 6. Threats to validity are discussed in Section 7. Section 8 provides the conclusion of the article.
@@ 3 Related Work
There are three recent peer-reviewed surveys about web testing that are close to our work [31, 32, 35], analyzing from 26 to 98 (but this latter going back to 2008) articles, considering a similar time period to our 2014-2025 (see Table 1), although not covering the last two years (i.e., 2024-2025). By analyzing $259$ articles, we can provide a more thorough view of the last decade in web testing research. Furthermore, with 14 research questions, we do analyze several important aspects of web testing that were not discussed before. For example, considering the current success stories of artificial intelligence techniques in many different aspects of science and engineering, we make sure to analyze how they have been used in web testing (as we will discuss in Section 5.13). Also, our survey focus only on testing. Other aspects like “automated repair” of code [36] and “fault localization” [37] (which were included for example in [35]) are not analyzed here in our survey, as not in scope.
A non-peer-reviewed survey [34], published on arXiv in 2024, covers similar topics compared to our survey. In [34] it is stated that 314 articles are reviewed, but only 199 are cited. So we cannot check if in our survey we missed any, or those other articles cover different aspects of web applications that we do not consider in our survey. That survey does not cover the latest 2 years, i.e., the time-period 2024-2025. Furthermore, with only six research questions [34], there are several important research questions that we study in our survey that are not covered in [34]. These for example include an analysis of the kinds of scientific contributions (RQ4), how studies with human subjects were performed (RQ6), how novel tools were compared (RQ9), which frameworks are most used to enable GUI-frontend testing (RQ12), the role of AI-techniques (RQ13) and details of applications in industry (RQ14).
Compared to existing surveys in the literature, our survey provides the important contribution of a more in-depth analysis of the last 12 years of research in web frontend testing.
@@ 4.1 Research questions
  • 1.
    RQ1: How many scientific articles were published each year between 2014-2025?
  • 2.
@@
  • 3.
    RQ3: What are the main scientific topics of focus in these articles?
  • 4.
    RQ4: What kind of scientific contributions do these articles provide?
  • 5.
@@
  • 6.
    RQ6: In human studies, what is the number of participants involved in the experiments?
  • 7.
    RQ7: Which tools are introduced or improved?
  • 8.
    RQ8: How often tools are open-source?
  • 9.
    RQ9: In algorithm experiments, how often are proposed tools compared with other tools?
  • 10.
    RQ10: In case of algorithm experiments, how many faults are found?
  • 11.
    RQ11: Which test generation technique is used most?
  • 12.
@@
RQ1 and RQ2 aim to provide a general overview of the demographics of the publications. RQ3 and RQ4 focus on identifying the main topics and contributions of the papers, acknowledging that Web Testing is a broad and diverse field with several areas of focus. In cases where the paper discusses an experiment, we discuss the used SUTs in RQ5. If the experiments involve human participants, RQ6 addresses this aspect. Through RQs 7,8 and 9, we aim to analyze the tools presented in the studies and understand their lifespan, usage and maintenance. RQ10 attempts to check the effectiveness of the proposed tools and techniques. RQs 11, 12 and 13 attempt to better understand the approaches on the proposed methods and techniques. RQ11 examines the balance between white-box and black-box testing approaches, and RQ12 investigates the extent to which libraries and frameworks are used to facilitate GUI testing. RQ13 and RQ14 address two particularly important aspects: RQ13 explores the integration of AI techniques in web testing, while RQ14 evaluates the degree of collaboration between academia and industry within the selected studies. We believe that these research questions address various key aspects and provide a comprehensive understanding of the current state of the literature.
@@ 4.2 Databases and Search Queries
IEEE ((”All Metadata”: ”web NEAR/2 test*”) OR (”All Metadata”: ”web NEAR/2 fuzz*”) OR (”All Metadata”: ”UI NEAR/2 test*”) OR (”All Metadata”: ”UI NEAR/2 fuzz*”) OR (”All Metadata”: ”GUI NEAR/2 test*”) OR (”All Metadata”: ”GUI NEAR/2 fuzz*”) OR ((”Document Title”: ”GUI”) AND ((”All Metadata”: ”test”) OR (”All Metadata”: testing)) AND (”All Metadata”: ”web”)) OR ((”Document Title”: test* OR ”Document Title”: fuzz*) AND (”Document Title”: ”web app*”))) AND NOT((”All Metadata”: fuzzy)) AND (NOT (”Document Title”: mobile OR ”Document Title”: android OR ”Document Title”: iOS OR ”Document Title”: desktop)) phase 1 - 160 phase 2 - 37
ACM Title: (test*) AND (NOT (Title:(mobile) OR Title: (android))) AND ((Abstract:(GUI) OR Abstract:(front-end) OR Abstract:(frontend) OR Abstract:(front end)) AND (Abstract:(test*) OR Abstract:(fuzz*)) AND (Abstract: (Web))) phase 1 - 58 phase 2 - 5
ScienceDirect Title, abstract, keywords: (”test” OR ”testing” OR ”fuzz” OR ”fuzzing”) AND (”user interface” OR ”UI” OR ”GUI”) AND (”web application” OR ”web”) phase 1 - 57 phase 2 - 32
Wiley ”test* OR fuzz*” in Abstract and ”GUI OR frontend OR front-end OR front end OR user interface OR UI” in Abstract and ”test*” in Title phase 1 - 13 phase 2 - 0
Web of Science (((TI=(”test*” OR ”testing*” OR ”fuzz*” OR ”fuzzing*”)) AND AB=(”test*” OR ”testing*” OR ”fuzz*” OR ”fuzzing*”)) AND AB=(”web”)) AND AB=(”GUI” OR ”graphical user interface” OR ”UI” OR ”user interface” OR ”front end” OR ”front-end” OR ”frontend”) phase 1 - 61 phase 2 - 6
SpringerLink (”test” OR ”testing” OR ”fuzz” OR ”fuzzing”) AND (”web application”) AND (”GUI” OR ”graphical user interface” OR ”UI” OR ”user interface” OR ”front end” OR ”front-end” OR ”frontend”) phase 1 - 236 phase 2 - 161
MIT Title containing these search terms: testing OR test OR fuzz OR fuzzing AND Abstract/description containing these search terms: ”test” OR ”testing” OR ”fuzz” OR ”fuzzing” AND Abstract/description containing these search terms: ”GUI” OR ”front end” OR ”front-end” OR ”UI” OR ”user interface” AND Abstract/description containing these search terms: ”web application” phase 1 - 0 phase 2 - 20
Total phase 1 - 625 phase 2 - 261
Table  LABEL:table:dbqueries shows the queries we run in every database.
@@ 4.3 Paper Selection Criteria
Our initial search, conducted in 17-24/04/2024 resulted in 625 papers. This large number can be attributed to the fact that our main keywords are quite common and appear even in papers that might not be related to web testing. We decided to eliminate the papers that did not fit the criteria below.
  • 1.
@@
The first two authors went over all the papers independently, decided on the list of papers to be included, and then compared their respective lists. In case of any discrepency, the issue was solved by the fourth author. After checking all the papers retrieved initally from the databases, we ended up with 145 papers. These papers were reviewed again from the authors, and another 21 papers were excluded.
The second phase in paper collection involves snowballing. We followed the guidelines described in [38]. We conducted forward and backward snowballing, and we ended up adding 90 new papers. Our final list amounts to 214 papers. Based on this selection of articles, a first draft of this survey was published on arXiv [39].
To get a better picture of the current state-of-the-art in web testing, in particular to take into account the latest development in AI-based testing, we went through a third phase during 12-15/05/2025, where we added and analyzed papers published in 2024 and 2025. We executed again the steps described in phases one and two, this time setting the publication years 2024 and 2025. After the third phase, the total number of papers included in this study is $259$.
Figure 1: Paper selection process
@@ 4.4 Data Extraction
The study comprises an extensive collection of papers. To manage and analyze the data, a structured Google spreadsheet was created. This document contains the list of selected papers alongside the extracted data, which is organized to address the research questions outlined in Section 4.1. Given the extensive collection of papers analyzed in this study, the dataset was divided between the first two authors for initial review and data extraction. Upon completion of the spreadsheet containing all extracted data, the results were collaboratively reviewed and thoroughly cross-verified by the remaining two authors to ensure accuracy and consistency.
@@ 5 Empirical Study
5.1 RQ1- How many scientific articles were published each year between 2014-2025?
@@
Figure 2 illustrates the publication rate of papers on web testing from January 2014 to May 2025. The publication rate has remained relatively stable, fluctuating between 15 and 31 papers per year, with the peak reached in 2023. In 2025 up to May, 15 papers have already been published, indicating a promising trend that the rate of publication is likely to continue without decline.
RQ1: The field of research on Web Testing has been significantly active in the last decade, with no significant signs of any slow down.
@@ 5.2 RQ2- Which are the most common venues where these articles are published?
Table 9 shows all the venues in which the different articles have been published. These venues are sorted based on the number of published articles. For each venue, it is indicated if it is a journal, a conference or a workshop. Due to the large number of different venues (more than 100), we only report the acronym of the venues in which at lest two articles have been published. All others are grouped together into the “Other conferences” and “Other journals” categories.
@@
Regarding journals, the most commons are JSS (Journal of Systems and Software), STVR (Software Testing, Verification and Reliability), SQJ (Software Quality Journal), TSE (IEEE Transactions on Software Engineering), EMSE (Empirical Software Engineering), IST (Information and Software Technology) and TOSEM (ACM Transactions on Software Engineering and Methodology). These are arguably the most important journals focused on software engineering and software testing.
@@
RQ2: ICST is the most common venue where to find the latest advances in Web Testing research, followed by FSE and ICSE. The top journals in software engineering all share a similar amount of articles on Web Testing.
5.3 RQ3- What are the main scientific topics of focus in these articles?
@@
Table 3: Categories and Total papers per category
Category Nr.Papers
Test Generation 124
Test Maintenance 52
Testing Framework 49
Model Inference 35
Oracle 21
Coverage Criteria 14
Test Prioritization 7
Taxonomy 4
Dataset 2